# Cmprsr: Abstractive Token-Level Question-Agnostic Prompt Compressor

### v1

@@ Abstract
Motivated by the high costs of using black-box Large Language Models (LLMs), we introduce a novel prompt compression paradigm, under which we use smaller LLMs to compress inputs for the larger ones. We present the first comprehensive LLM-as-a-compressor benchmark spanning $25$ open- and closed-source models, which reveals significant disparity in models’ compression ability in terms of (i) preserving semantically important information (ii) following the user-provided compression rate (CR). We further improve the performance of gpt-4.1-mini, the best overall vanilla compressor, with Textgrad-based compression meta-prompt optimization. We also identify the most promising open-source vanilla LLM—Qwen3-4B—and post-train it with a combination of supervised fine-tuning (SFT) and Group Relative Policy Optimization (GRPO), pursuing the dual objective of CR adherence and maximizing the downstream task performance. We call the resulting model Cmprsr and demonstrate its superiority over both extractive and vanilla abstractive compression across the entire range of compression rates on lengthy inputs from MeetingBank and LongBench as well as short prompts from GSM8k. The latter highlights Cmprsr’s generalizability across varying input lengths and domains. Moreover, Cmprsr closely follows the requested compression rate, offering fine control over the cost-quality trade-off.
Figure 1: Extractive compression selects a subset of the input sequence tokens, while abstractive compression allows for clever paraphrases. While one can use vanilla LLMs as abstractive compressors, their performance can be further improved with RL-based post-training, yielding Cmprsr. Note that extractive compression may introduce ambiguities, e.g. “February April”, “mission forward 50 years” bbc2025, while vanilla abstractive compression does not adhere to the desired compression rate.
@@ 1 Introduction
The discovery of scaling laws kaplan2020scalinglawsneurallanguage set the trend for training increasingly large Language Models (LMs). Despite rapid advancements in both hardware and software supporting LLMs’ inference, the costs of their usage continue to surge. This trend reflects not only growing adoption liang2025widespreadadoptionlargelanguage, but also the Jevons paradox Luccioni25: efficiency gains that spur even greater consumption.
According to recent estimates tully2025midyear, $87\%$ of the company’s spending on LLMs is attributed to black-box LLMs, accessed via API. This means that in practice the only way the majority of LLM consumers can optimize their spendings is through minimizing the queries’ length passed to the models; this can be achieved with the token-level prompt compression li-etal-2025-prompt, i.e. exploiting the redundancy of human language Shannon51 and mapping the original sequence to a shorter one, while preserving original semantics.
We focus on the question-agnostic compression jiang-etal-2023-llmlingua, which aims to process the context provided to the model so that the compression can be re-used for any context-related question or task: possible use-cases include (i) combining compression with Retrieval-Augmented Generation Gao2023RetrievalAugmentedGF via retrieving pre-compressed entries, (ii) single-call compressions of lengthy texts such as meeting transcripts hu-etal-2023-meetingbank for their subsequent comprehensive analysis involving multiple LLM calls, (iii) optimizing LLM-powered learning platforms via compressing learning materials (iv) addressing tokens-per-minute (TPM) API bandwidth limitations.
The most popular approaches tailored for this set-up are extractive jiang-etal-2023-llmlingua; pan-etal-2024-llmlingua, meaning that compression is posed as a binary classification problem of preserving/removing each of the input sequence parts (usually tokens). We hypothesize that abstractive methods, operating under a much larger space of valid compressions, can provide better outputs through paraphrase of the input sequence. To this end, we define $2$ metrics, characterizing the quality of a Compressor LLM: (i) Target LLM performance on a downstream task given its inputs are preprocessed by the Compressor (ii) Adherence of the Compressor to the target Compression Rate (CR), reflecting user’s tolerance to the quality deterioration versus the incurred costs. We ask the following research questions (RQs):
@@
  • RQ2. Can we further improve their performance with prompt optimization techniques such as yuksekgonul2025optimizing?
@@
We share results of the extensive benchmarking addressing RQ1, and show that the answer to RQ2 is positive in case the prompt is used with the same Compressor model it was optimized for. Most importantly, we present Cmprsr, an abstractive Compressor LLM that outperforms SOTA extractive compression across different CR-s, which answers RQ3.
@@ 2 Related work
Compression can be performed in either question-aware shandilya2024tacorltaskawareprompt; kim2025acornnoiserobustabstractivecompression; choi-etal-2024-reading; yoon-etal-2024-compact or question-agnostic way. In the question-aware set-up, compression is conditioned on the given question and aims to filter out all irrelevant information. While allowing high CRs, this also prevents compression re-usage for the new queries. Motivated by the use-cases detailed in Sec. 1, we focus on the question-agnostic set-up.
Furthermore, prompt compression techniques is divided into $2$ categories li-etal-2025-prompt:
@@
Embedding-level (a.k.a soft prompt methods mu2023learning; chevalier-etal-2023-adapting), which require access to the internals of the Target model, and are therefore not viable for the black-box models.
@@
Token-level (a.k.a hard prompt methods): (i) Extractive: compression space is the set of order-preserving subsequences of the input; the filtering can be implemented on the token jiang-etal-2023-llmlingua; pan-etal-2024-llmlingua or the sentence Liskavets25 level. Notably, hu2025dynamiccompressingpromptsefficient; shandilya2024tacorltaskawareprompt use RL techniques to optimize the extraction policy. (ii) Abstractive: the compression space is all sequences over the vocabulary, which allows semantics-preserving paraphrase/reordering via tokens not present in the input.
@@
Below, we detail prior contributions falling into the same broad category as ours, i.e. question-agnostic token-level abstractive compression. pu-etal-2024-style rely on vanilla LLaMA-2-7B to perform compression guided by demonstrations, optimized for the particular dataset. Despite improvement over straightforward prompting, this method is not readily generalizable across tasks, as each new dataset requires generating/selecting a new set of demonstrations. In a recent work, zhang2025scopegenerativeapproachllm develop another approach to abstractive compression, using either GPT-4o-mini or Qwen-2.5: they rely on dynamic chunking and chunk-specific CRs, which allows to preserve important context.
@@
Neither of the two above-mentioned methods compare a meaningful number of LLMs in terms of their aptitude for compression, which, as we show, greatly varies. Furthermore, according to our experiments, even the models dominating the ”vanilla compression leaderboard” can be significantly improved with post-training in terms of both compression quality and adherence to the user-specified CR. The work of chuang-etal-2024-learning most closely addresses this limitation: while it still experiments with a single backbone (Vicuna-7B), the model is actually tuned for better compressions. The main limitation of the presented approach is that it falls short of adopting RL: Vicuna-7B serves as both Compressor and the Target model, and the signal comes from the semantic preservation loss between the original and the compressed input activations.
Among the works targeting specific downstream tasks, larionov-eger-2025-promptoptme investigate compression for the machine translation quality assessment. While the resulting model cannot be used for prompt compression ”in the wild”, the authors notably introduce RL —in the form of preference optimization (ORPO)—for training an abstractive compressor.
@@ 3 CompressionBench
Figure 2 illustrates the 2 evaluation pipelines used to assess the impact of compression on model utility: summarization and question answering (QA) pipelines. The template of the Compressor prompt is provided in the Appendix. It includes length conditioning, where the desired CR is converted into a target token count and appended to the system prompt. For both the benchmarking and the training of Cmprsr, transcripts are segmented into manageable chunks before being passed to the Compressor model, following the methodology of pan-etal-2024-llmlingua to avoid truncation of incomplete sentences. Compressed chunks are then concatenated to reconstruct the full compressed transcript. We evaluate the resulting compressions through two pipelines:
Summarization pipeline. The goal of this pipeline is to evaluate how well the compressed transcripts preserve the semantic content necessary for high-quality summaries. We generate summaries for both the original and the compressed transcripts using the same Target model, and compute the BERTScore-F1 between them. This measures semantic overlap and serves as an intuitive proxy for information retention: higher BERTScore indicates that the compressed input preserves the key content of the original.
Figure 2: MeetingBank evaluation pipeline. We assess transcripts’ compressions on 22 downstream tasks. (i) Summarization, i.e we compute BertScore between compressed and original. (ii) QA, where we build a dataset of questions and answers from MeetingBank transcripts, and measure the Target model’s accuracy using the compressed context.
Table 1: Compression performance of various vanilla models on the MeetingBank transcripts: truncated version of Table 4. Within each group, we sort the models based on the average QA performance accross compression rates.
@@
    Vanilla LLMs poorly adhere to the prompted CR. While they are susceptible to the prompted rate, the length of the generations skews towards some fixed CR, leading to ”under-compression” for high CRs (0.1) and ”over-compression” for low CRs (0.5). LLMLingua-2 does not suffer from this limitation, as the classification threshold can be dynamically adjusted for the extractive methods.
@@
The results from this section inform our choice of the models for the Cmprsr experiments: Qwen3-4B among the small open-source models (it performs on par with LLams-3.2-3B with slighly better CR adherence), and gpt-4.1-mini among the closed-source ones.
@@ 3.1 Boosting vanilla performance with TextGrad
Motivation. The system prompt is crucial to aligning LLM outputs with user expectations. Therefore, it strongly affects the performance of our LLM-based compressor. As LLMs have grown in popularity, prompting has matured, offering diverse techniques such as step-by-step reasoning and few-shot examples. Yet selecting the right technique and phrasing remains nontrivial and task-dependent. To address this, we employ TextGrad yuksekgonul2024textgradautomaticdifferentiationtext for principled prompt optimization. TextGrad treats texts as optimization variables, refining them via natural-language “gradients.”
Method. We model the MeetingBank QA benchmark as a computation graph whose nodes represent LLM pipeline stages. Each node has a role description to help the optimizer-LLM grasp the high-level dependencies. Beginning with ground-truth evaluation at the output, TextGrad (i) identifies failures, (ii) proposes fixes, (iii) backpropagates feedback through the graph, and (iv) updates the compressor’s system prompt. We evaluate QA performance using a 10-point LLM judge that grades each answer. We also assess adherence to the target CR. Further details appear in Appendix C. We applied TextGrad to optimize the system prompt for two well-performing compressor LLMs, gpt-4.1-mini and Qwen3-4B. The optimization trajectories differed significantly between the two models.
TextGrad for Qwen3-4B: Figure 3 shows QA quality ($y$-axis) versus compression-rate adherence ($x$-axis) on a 100-transcript hold-out set, with each point representing a TextGrad iteration. Panels display parts of the initial (iter. 0) and best (iter. 8) prompts.
The plot reveals a trade-off: longer outputs (weaker adherence) retain more relevant information, boosting the LLM-Judge score. The optimization trajectory oscillates, with TextGrad alternating between improving quality and tightening adherence. The resulting prompts form a Pareto front, enabling choice between strict CR adherence and slightly longer, higher-quality compressions. The best prompt yields a 0.51 gain in the average LLM-Judge score with only a 0.02 drop in CR adherence. It addresses failure modes by adding positive and negative examples and emphasizing named entities, steering the compressor to retain key information while staying within budget.
Figure 3: TextGrad system prompt optimization for Qwen3-4B.
@@
Figure 4: TextGrad prompt optimization for gpt-4.1-mini.
@@ 4.1 Supervised fine-tuning (SFT)
This stage has two goals: (i) improving Qwen3-4B performance by distilling gpt-4.1-mini compression strategies, and (ii) enhancing adherence to the prompted CR. To this end, we append the true token count to the prompt, i.e., the fine-tuning sequence is Original + Length Conditioned Prompt (len(Compression)) + Compression, and the loss is computed on the Compression part; this “length conditioning” is inspired by Hindsight Instruction Relabelling (HIR) Zhang23; pie_iclr_2024_spotlight. Preliminary experiments also tested “quality conditioning,” i.e., passing the normalized compression quality relative to others in the same rate range, but it had limited success. We varied the learning rate in $10^{-4},10^{-8}]$ and selected the SFT checkpoint with the best validation accuracy: the one trained on the upsampled dataset with $lr=10^{-5}$.
@@ 4.2 GRPO
Training. Further details of GRPO training are given in Sec. D.1. While we also experimented with DPO and ORPO, these approaches did not outperform the GRPO-based pipeline; we describe these setups in Sec. D.2 of the Appendix.
@@ 4.3 Evaluation
To assess the effectiveness of our approach, we evaluate model performance under two complementary task settings as explained in Section 3: summarization and QA. Our primary focus is (i) to examine whether the compressed representations are question-agnostic, i.e., they preserve general information that supports diverse queries, and (ii) to verify whether they retain sufficient abstraction to follow the intent of summarization tasks. Although the model is trained exclusively on the MeetingBank dataset, we evaluate it across multiple out-of-distribution (OOD) benchmarks, including GSM8K for QA and LongBench for both summarization and QA. This setup tests generalization across tasks with differing requirements, ranging from factual short reasoning to long-document summarization. The evaluation pipeline operates as follows: a compressed text $x_{C}$ is first generated, which is then fed into the solver model to either produce a summary or answer task-specific questions.
@@
We report the evaluation results in Table 2. For summarization tasks, we adopt the Bert-F1 score as the primary metric, while for QA tasks we use accuracy (i.e 1 if the predicted answer matches the ground truth, and 0 otherwise). For LongBench, which includes multiple subtasks within the same dataset, we aggregate results by averaging over all QA-related tasks (e.g., single-document QA and few-shot QA) and separately averaging over all summarization-related tasks. Full, fine-grained results for each subtask are provided in the Appendix E.
Table 2: Model performance across QA and summarization benchmarks under varying CRs.
CR Model Question Score Summary Score GSM8K MeetingBank LongBench MeetingBank LongBench $\Delta_{CR}$ EM $\Delta_{CR}$ QA Score $\Delta_{CR}$ QA Score $\Delta_{CR}$ BERTF1 $\Delta_{CR}$ BERTF1 0.1 Lingua-2 0.02 0.01 -0.01 0.17 -0.01 0.17 -0.01 0.864 -0.01 0.856 GPT-4.1-mini 0.22 0.50 0.03 0.16 0.12 0.21 0.03 0.882 0.06 0.876 Qwen3-4B 0.36 0.67 0.07 0.17 0.12 0.19 0.07 0.877 0.11 0.873 TextGrad-Qwen 0.39 0.59 0.08 0.20 0.12 0.21 0.08 0.884 0.10 0.877 SFT-Qwen 0.30 0.59 0.04 0.15 0.06 0.19 0.04 0.876 0.05 0.866 Cmprsr 0.03 0.13 0.00 0.20 0.00 0.18 0.00 0.878 0.01 0.871 0.3 Lingua-2 0.02 0.13 -0.03 0.33 -0.01 0.20 -0.03 0.889 -0.02 0.877 GPT-4.1-mini 0.20 0.83 -0.01 0.26 0.15 0.21 -0.01 0.895 0.09 0.882 Qwen3-4B 0.21 0.74 -0.06 0.20 0.02 0.19 -0.06 0.887 -0.01 0.875 TextGrad-Qwen 0.25 0.76 -0.02 0.25 0.06 0.20 -0.02 0.894 0.01 0.882 SFT-Qwen 0.21 0.79 -0.04 0.23 0.02 0.21 -0.04 0.892 -0.01 0.874 Cmprsr 0.03 0.46 0.02 0.35 0.04 0.19 0.03 0.903 0.01 0.886 0.5 Lingua-2 0.00 0.54 -0.03 0.41 -0.01 0.16 -0.03 0.905 -0.02 0.888 GPT-4.1-mini 0.08 0.89 -0.13 0.29 0.04 0.22 -0.13 0.900 -0.02 0.884 Qwen3-4B 0.06 0.79 -0.21 0.22 -0.13 0.20 -0.21 0.889 -0.17 0.878 TextGrad-Qwen 0.19 0.78 -0.14 0.28 -0.05 0.22 -0.14 0.895 -0.12 0.882 SFT-Qwen 0.06 0.84 -0.18 0.28 -0.10 0.23 -0.18 0.894 -0.14 0.875 Cmprsr -0.03 0.72 0.03 0.43 0.02 0.18 0.03 0.911 0.01 0.895
Based on Table 2, higher raw scores are often achieved at the cost of exceeding the target CR, i.e., resulting in positive $\Delta_{CR}$. Since our objective is to maximize task performance under a relaxed constraint of $\Delta_{CR}\leq\epsilon$, we do not simply select the highest-scoring models. Instead, we focus on models that respect the CR budget while still delivering competitive outcomes. For this study, we set $\epsilon$ to 0.03. Note that perfect adherence to the CR target is often unattainable, and slightly negative $\Delta_{CR}$ is preferable.
Under this criterion, SFT and TextGrad occasionally yield strong performance, but only when higher $\Delta_{CR}$ values are tolerated, meaning the models compress less than required. To address this, we explore GRPO on top of SFT. This two-stage approach, called Cmprsr, uses a raw-stream version for SFT while applying a distilled version for GRPO. The framework allows us to combine task-specific gradients with compression-aware reinforcement in a second refinement stage. Our proposed Cmprsr model consistently achieves the best trade-off across both QA and summarization tasks. By effectively combining SFT with GRPO, it maintains $\Delta_{CR}\leq 0$, ensuring strict adherence to the target CR while outperforming baseline models. Figure 5 further illustrates model adherence to the CR across training stages.
@@
4.4 Practical applicability discussion
While Cmprsr outperforms LLMLingua-2 across the CR spectrum, it is important to note that (i) it is based on a 4B backbone, while LLMLingua-2 is a fine-tuned xlm-roberta-large of 0.55B parameters (ii) the autoregressive nature of Cmprsr implies latency overhead. This, however, is unlikely to prevent widespread adoption of LLM-based compressors. First, most of the question-agnostic use-cases discussed in Sec. 1 allow for the the usage of the pre-computed compressions, which decreases the importance of latency. Secondly, the usage of either 0.55B or a 4B model will come at a fraction of the savings, generated by optimized inference of the Target model. Quantitatively, LLama-3.2-3B inference cost is $\mathdollar 0.06$ (both input and output at together.ai) and GPT-5 is $\mathdollar 1.25$ per $10^{6}$ input tokens, meaning that in case of $CR=0.3$, the input-incurred GPT-5 cost would be reduced by $64\%$, while in the limit of the free compression the savings would amount to $70\%$.

### v2

@@ Abstract
Motivated by the high costs of using black-box Large Language Models (LLMs), we introduce a novel prompt compression paradigm, under which we use smaller LLMs to compress inputs for the larger ones. We present the first comprehensive LLM-as-a-compressor benchmark spanning $25$ open- and closed-source models, which reveals significant disparity in models’ compression ability in terms of (i) preserving semantically important information (ii) following the user-provided compression rate (CR). We further improve the performance of gpt-4.1-mini, the best overall vanilla compressor, with Textgrad-based compression meta-prompt optimization. We also identify the most promising open-source vanilla LLM—Qwen3-4B—and post-train it with a combination of supervised fine-tuning (SFT) and Group Relative Policy Optimization (GRPO), pursuing the dual objective of CR adherence and maximizing the downstream task performance. We call the resulting model Cmprsr and demonstrate its superiority over both extractive and vanilla abstractive compression across the entire range of compression rates on lengthy inputs from MeetingBank and LongBench as well as short prompts from GSM8k. The latter highlights Cmprsr’s generalizability across varying input lengths and domains. Moreover, Cmprsr closely follows the requested compression rate, offering fine control over the cost-quality trade-off. Importantly, we show that in case of lengthy inputs, compression can be practically lossless in terms of the downstream task performance.
Cmprsr: Abstractive Token-Level Question-Agnostic Prompt Compressor
Ivan Zakazov1,2   Berke Argın1,2thanks: Authors contributed equally.   Oussama Gabouj1,2footnotemark:   Kamel Charaf1,2footnotemark: Alexander Sharipov2footnotemark:   Alexi Semiz2   Lorenzo Drudi2   Nicolas Baldwin2 Robert West2 1Compresr Inc.   2EPFL {ivan, berke, oussama, kamel}@compresr.ai
{alexander.sharipov, robert.west}@epfl.ch
@@ 1 Introduction
The discovery of scaling laws Kaplan et al. (2020) set the trend for training increasingly large Language Models (LMs). Despite rapid advancements in both hardware and software supporting LLMs’ inference, the costs of their usage continue to surge. This trend reflects not only growing adoption Liang et al. (2025), but also the Jevons paradox Luccioni et al. (2025): efficiency gains that spur even greater consumption.
According to recent estimates Tully et al. (2025), $87\%$ of the company’s spending on LLMs is attributed to black-box LLMs, accessed via API. This means that in practice the only way the majority of LLM consumers can optimize their spendings is through minimizing the queries’ length passed to the models; this can be achieved with the token-level prompt compression Li et al. (2025), i.e. exploiting the redundancy of human language Shannon (1951) and mapping the original sequence to a shorter one, while preserving original semantics.
We focus on the question-agnostic compression Jiang et al. (2023), which aims to process the context provided to the model so that the compression can be re-used for any context-related question or task: possible use-cases include (i) combining compression with Retrieval-Augmented Generation Gao et al. (2023) via retrieving pre-compressed entries, (ii) single-call compressions of lengthy texts such as meeting transcripts Hu et al. (2023) for their subsequent comprehensive analysis involving multiple LLM calls, (iii) optimizing LLM-powered learning platforms via compressing learning materials (iv) addressing tokens-per-minute (TPM) API bandwidth limitations.
The most popular approaches tailored for this set-up are extractive Jiang et al. (2023); Pan et al. (2024), meaning that compression is posed as a binary classification problem of preserving/removing each of the input sequence parts (usually tokens). We hypothesize that abstractive methods, operating under a much larger space of valid compressions, can provide better outputs through paraphrase of the input sequence. To this end, we define $2$ metrics, characterizing the quality of a Compressor LLM: (i) Target LLM performance on a downstream task given its inputs are preprocessed by the Compressor (ii) Adherence of the Compressor to the target Compression Rate (CR), reflecting user’s tolerance to the quality deterioration versus the incurred costs. We ask the following research questions (RQs):
@@
  • RQ2. Can we further improve their performance with prompt optimization techniques such as Yuksekgonul et al. (2025)?
@@
We share results of the extensive benchmarking addressing RQ1, and show that the answer to RQ2 is positive in case the prompt is used with the same Compressor model it was optimized for. Most importantly, we present Cmprsr, an abstractive Compressor outperforming SOTA extractive compression across different CR-s, which answers RQ3.
@@ 2 Related work
Figure 2: Evaluation pipeline. We assess transcripts’ compressions on $2$ downstream tasks. (i) Summarization, i.e we compute BertScore between compressed and original. (ii) QA, where we build a dataset of questions and answers from MeetingBank transcripts, and measure the Target model’s accuracy using the compressed context.
@@
Compression can be performed in either question-aware Shandilya et al. (2024); Kim et al. (2025); Choi et al. (2024); Yoon et al. (2024) or question-agnostic way. In the question-aware set-up, compression is conditioned on the given question and aims to filter out all irrelevant information. While allowing high CRs, this also prevents compression re-usage for the new queries. Motivated by the use-cases detailed in Sec. 1, we focus on question-agnostic set-up. According to Li et al. (2025), prompt compression can also be divided into the following $2$ categories:
@@
Embedding-level (a.k.a soft prompt methods Mu et al. (2023); Chevalier et al. (2023)), which require access to the Target model, and are therefore not viable for the black-box models.
@@
Token-level (a.k.a hard prompt methods): (i) Extractive: compression space is the set of order-preserving subsequences of the input; the filtering can be implemented on the token Jiang et al. (2023); Pan et al. (2024); Zhao et al. (2025) or the sentence Liskavets et al. (2025) level. Notably, Hu et al. (2025); Shandilya et al. (2024) use RL techniques to optimize the extraction policy. Filtering can be informed by information entropy Jiang et al. (2023); Li et al. (2023), attention scores Zhao et al. (2025) or performed with a classification model Pan et al. (2024) (ii) Abstractive: the compression space is all sequences over the vocabulary, which allows semantics-preserving reordering via tokens not present in the input.
@@
Below, we detail prior contributions falling into the same broad category as ours, i.e. question-agnostic token-level abstractive compression. Pu et al. (2024) rely on vanilla LLaMA-2-7B to perform compression guided by demonstrations, optimized for the particular dataset. Despite improvement over straightforward prompting, this method is not readily generalizable across tasks, as each new dataset requires generating/selecting a new set of demonstrations. In a recent work, Zhang et al. (2025) develop another approach to abstractive compression, using either GPT-4o-mini or Qwen-2.5: they rely on dynamic chunking and chunk-specific CRs, which allows to preserve important context.
@@
Neither of the two above-mentioned methods compare a meaningful number of LLMs in terms of their aptitude for compression, which, as we show, greatly varies. Furthermore, according to our experiments, even the models dominating the "vanilla compression leaderboard" can be significantly improved with post-training in terms of both compression quality and adherence to the user-specified CR. The work of Chuang et al. (2024) most closely addresses this limitation: while it still experiments with a single backbone (Vicuna-7B), the model is actually tuned for better compressions. The main limitation of the presented approach is that it falls short of adopting RL: Vicuna-7B serves as both Compressor and the Target model, and the signal comes from the semantic preservation loss between the original and the compressed input activations. Among the works targeting specific downstream tasks, Larionov and Eger (2025) investigate compression for the machine translation quality assessment. While the resulting model cannot be used for prompt compression "in the wild", the authors notably introduce RL (ORPO) for training an abstractive compressor.
@@ 3 CompressionBench
Figure 2 illustrates the 2 evaluation pipelines used to assess the impact of compression on model utility: summarization and question answering (QA) pipelines. The template of the Compressor prompt is provided in the Appendix. It includes length conditioning, where the desired CR is converted into a target token count and appended to the system prompt. For both the benchmarking and the training of Cmprsr, transcripts are segmented into manageable chunks before being passed to the Compressor model, following the methodology of Pan et al. (2024) to avoid truncation of incomplete sentences. Compressed chunks are then concatenated to reconstruct the full compressed transcript. We evaluate the resulting compressions through two pipelines:
Table 1: Compression performance of various vanilla models on the MeetingBank transcripts; truncated version of Table 4. Within each group, we sort the models based on the average QA performance across compression rates.
@@
Summarization pipeline. The goal of this pipeline is to evaluate how well the compressed transcripts preserve the semantic content necessary for high-quality summaries. We generate summaries for both the original and the compressed transcripts using the same Target model, and compute the BERTScore-F1 between them: a higher BERTScore indicates that the compressed input preserves the key content of the original.
@@
    Vanilla LLMs poorly adhere to the prompted CR. While they are susceptible to the prompted rate, the length of the generations skews towards some fixed CR, leading to "under-compression" for high CRs (0.1) and "over-compression" for low CRs (0.5). LLMLingua-2 does not suffer from this limitation, as the classification threshold is dynamically adjusted for the extractive methods.
@@
The results from this section inform our choice of the models for the Cmprsr experiments: Qwen3-4B among the small open-source models (it performs on par with LLama-3.2-3B with slighly better CR adherence), and gpt-4.1-mini among the closed-source ones.
@@ 3.1 Boosting vanilla performance with TextGrad
Motivation. As LLMs have grown in popularity, prompting has matured, offering diverse techniques such as step-by-step reasoning and few-shot examples. Yet selecting the right technique and phrasing remains nontrivial and task-dependent. To address this, we employ TextGrad Yuksekgonul et al. (2024) for principled prompt optimization. TextGrad treats texts as optimization variables, refining them via natural-language “gradients.”
Method. We model the MeetingBank QA benchmark as a computation graph whose nodes represent LLM pipeline stages. Each node has a role description to help the optimizer-LLM grasp the high-level dependencies. Beginning with ground-truth evaluation at the output, TextGrad (i) identifies failures, (ii) proposes fixes, (iii) backpropagates feedback through the graph, and (iv) updates the compressor’s system prompt. We evaluate QA performance using a 10-point LLM judge that grades each answer. We also assess adherence to the target CR. Further details appear in Appendix E. We applied TextGrad to optimize the system prompt for two well-performing compressor LLMs, gpt-4.1-mini and Qwen3-4B. The optimization trajectories differed significantly between the two models.
TextGrad for Qwen3-4B: Figure 3 shows QA quality ($y$-axis) versus compression-rate adherence ($x$-axis) on a 100-transcript hold-out set, with each point representing a TextGrad iteration. Panels display parts of the initial and best prompts.
The plot reveals a trade-off: longer outputs (weaker adherence) retain more relevant information, boosting the LLM-Judge score. The optimization trajectory oscillates, with TextGrad alternating between improving quality and tightening adherence. The resulting prompts form a Pareto front, enabling choice between strict CR adherence and slightly longer, higher-quality compressions. The best prompt yields a 0.51 gain in the average LLM-Judge score with only a 0.02 drop in CR adherence.
Figure 3: Prompt optimization for Qwen3-4B.
@@
Figure 4: Prompt optimization for gpt-4.1-mini.
@@ 4.1 Supervised fine-tuning (SFT)
This stage has two goals: (i) improving Qwen3-4B performance by distilling gpt-4.1-mini compression strategies, and (ii) enhancing adherence to the prompted CR. To this end, we append the true token count to the prompt, i.e., the fine-tuning sequence is Original + Length Conditioned Prompt (len(Compression)) + Compression, and the loss is computed on the Compression part; this “length conditioning” is inspired by Hindsight Instruction Relabelling (HIR) Zhang et al. (2023); Shypula et al. (2024). Preliminary experiments also tested “quality conditioning,” i.e., passing the normalized compression quality relative to others in the same rate range, but it had limited success. We varied the learning rate in $10^{-4},10^{-8}]$ and selected the SFT checkpoint with the best validation accuracy: the one trained on the upsampled dataset with $lr=10^{-5}$.
@@ 4.2 GRPO
Training. Further details of GRPO training are given in Sec. B.1. While we also experimented with DPO and ORPO, these approaches did not outperform the GRPO-based pipeline; we describe these setups in Sec. B.2 of the Appendix.
Figure 5: Performance comparison across QA and summary benchmarks. Left panel shows precision scores for QA benchmarks (GSM8K, Ruler, LongBench QA, InfiniteBench, LongBench v2, and MeetingBank QA). Right panel shows BERTScore F1 scores for summary benchmarks (MeetingBank Summary and LongBench Summary). The horizontal dashed lines indicate discrepancy between prompted and real compression ratios, and the horizontal black line represents the no-compression baseline. Detailed numerical results for compression rates and performance scores are also provided in Appendix C.
@@ 4.3 Evaluation
To assess the effectiveness of our approach, we evaluate model performance under two complementary task settings as explained in Section 3: summarization and QA. Our primary focus is (i) to examine whether the compressed representations are question-agnostic, i.e., they preserve general information that supports diverse queries, and (ii) to verify whether they retain sufficient abstraction to follow the intent of summarization tasks. Although the model is trained exclusively on the MeetingBank dataset, we evaluate it across multiple out-of-distribution (OOD) benchmarks, to test generalization. These include GSM8K for mathematical reasoning, MeetingBank (QA and Summarization), both version of LongBench, InfiniteBench, and Ruler. The evaluation pipeline operates as follows: a compressed text $x_{C}$ is first generated, which is then fed into the solver model (gpt-5-mini) to produce a summary / answer task-specific questions.
We compare Cmprsr with $2$ SOTA extractive methods – LLMLingua-2 and DAC Pan et al. (2024); Zhao et al. (2025) – and vanilla abstractive compression. We drop the original LLMLingua Jiang et al. (2023) and Selective context Li et al. (2023), as they were decisively outperformed by the baselines mentioned above. To achieve a more accurate comparison, we use the same backbone – Qwen-3-4B – for Cmprsr and DAC.
@@
We report the evaluation results in Figure 5 and Table 8 in the Appendix. For summarization tasks, we adopt the Bert-F1 score as the primary metric, while for QA tasks we use precision. For LongBench, LongBenchv2, InfiniteBench and Ruler, which include multiple subtasks within the same dataset, we aggregate results by averaging over QA-related tasks and separately averaging over summarization-related tasks. Full results for each subtask are provided in the Appendix C. Notably, DAC performance on GSM8k differs significantly from the original paper, as we run it in the zero-shot mode, and compress the question itself, while Zhao et al. (2025) compress demonstrations.
Based on Table 8, higher raw scores are often achieved at the cost of exceeding the target CR, i.e., resulting in positive $\Delta_{CR}$. To account for that, we determine the best method under the constraint of $\Delta_{CR}\leq\epsilon$. Under this criterion, Cmprsr consistently achieves the best trade-off across both QA and summarization tasks. By effectively combining SFT with GRPO, it maintains $\Delta_{CR}\leq 0$, ensuring strict adherence to the target CR while outperforming baseline models. Figure 6 further illustrates model adherence to the prompted CR across the training stages.
@@
Importantly, Figure 5 shows that the datasets with the longest entries (LongBench V2, Infinite bench) allow for almost lossless compression in terms of the downstream task performance.